# Quantum reservoir probing of quantum phase transitions

### v2

@@ METHODS
To construct the initial state, we define $\mathcal{H}_{\mathrm{DMRG}}$ that yields $|\Psi_{k}^{\mathrm{in}}\rangle$ as its ground state obtained from the density matrix renormalization group (DMRG) method [65, 66, 67]. For the transverse-field Ising model, the ANNNI model, and the spin-$1$ XXZ model, the initial states are prepared in product states as described in Eqs. (6) and (9). The Hamiltonian $\mathcal{H}_{\mathrm{DMRG}}$ is correspondingly defined, consisting of the local up or down magnetic field along the $z$-axis, which yields the all-up or Néel states as the ground state. A local magnetic field is applied to the central spin to prepare it as $|\psi(s_{k})\rangle$ in the spin-$1/2$ systems and $|\psi^{\prime}(s_{k})\rangle$ in the spin-$1$ system. For the eSSH model, we set $\mathcal{H}_{\mathrm{DMRG}}=\mathcal{H}(J^{\prime}=6.0)$, applying the $s_{k}$-dependent central field and an edge field in $+z$-axis with a strength of $0.1$. The initial state is then obtained as the ground state of $\mathcal{H}_{\mathrm{DMRG}}$ using the DMRG method, expressed as a matrix product state with a bond dimension $\chi$. For the time-evolution $e^{-i\mathcal{H}t}|\Psi_{k}^{\mathrm{in}}\rangle$, we employ the time-dependent variational principle [68, 69, 70] with a bond dimension $\chi$ and a time step $\Delta t=0.005$. The tensor network calculations are implemented using the Julia version of the ITensor library [71, 72].
@@
From a uniform distribution $[0,1]$, a set of $256$ input values $\{s_{k}\}$ is sampled. Each input $s_{k}$ produces a distinct initial state $|\Psi_{k}^{\mathrm{in}}\rangle$, and the dynamics of operator $\langle O_{i}(t)\rangle_{k}$ is subsequently calculated. Among them, $l^{\mathrm{tr}}=128$ instances are designated for training, while the remaining $l^{\mathrm{ts}}=128$ instances are reserved for testing. The linear transformation is represented using a 2D vector $\bm{x}_{i,k}(t)=\left(\langle O_{i}(t)\rangle_{k},1\right)$ and expressed as $y_{i,k}(t)=\bm{x}_{i,k}(t)\bm{w}_{i}(t)$, with the weight vector $\bm{w}_{i}(t)=(w_{i,O}(t),w_{i,c}(t))^{\top}$.
In the training phase, an ($l^{\mathrm{tr}}\times 2$)-dimensional matrix $X_{i}^{\mathrm{tr}}(t)=\{\bm{x}_{i,k}(t)\}_{k=1}^{l^{\mathrm{tr}}}$ is constructed by summing up the observed dynamics. The output, represented as an $l^{\mathrm{tr}}$-dimensional vector $\bm{y}_{i}^{\mathrm{tr}}(t)=\{y_{i,k}\}_{k=1}^{l^{\mathrm{tr}}}$, is calculated as $\bm{y}_{i}^{\mathrm{tr}}(t)=X^{\mathrm{tr}}_{i}(t)\bm{w}_{i}(t)$. The weight vector is trained to closely match the output $\bm{y}_{i}^{\mathrm{tr}}(t)$ with the original input ${\bm{{s}}}^{\mathrm{tr}}=\{s_{k}\}_{k=1}^{l^{\mathrm{tr}}}$. The optimal solution, which minimizes the least squared error between these vectors, is given by
$\bm{w}_{i}(t)={{X}_{i}^{\mathrm{tr}}}(t)^{+}{\bm{{s}}}^{\mathrm{tr}},$ (11)
where ${{X}_{i}^{\mathrm{tr}}}(t)^{+}$ denotes the Moore-Penrose pseudoinverse-matrix of ${{X}_{i}^{\mathrm{tr}}}(t)$.
In the testing phase, an ($l^{\mathrm{ts}}\times 2$)-dimensional matrix $X_{i}^{\mathrm{ts}}(t)=\{\bm{x}_{i,k}(t)\}_{k=l^{\mathrm{tr}}+1}^{l^{\mathrm{tr}}+l^{\mathrm{ts}}}$ is similarly composed, yielding the output $\bm{y}_{i}^{\mathrm{ts}}(t)=X_{i}^{\mathrm{ts}}(t)\bm{w}_{i}(t)$ using the trained weights. The estimation performance is assessed by comparing this testing output with the corresponding input ${\bm{{s}}}^{\mathrm{ts}}=\{s_{k}\}_{k=l^{\mathrm{tr}}+1}^{l^{\mathrm{tr}}+l^{\mathrm{ts}}}$ using the determination coefficient
$R^{2}_{i}(t)=\frac{\mathrm{cov}^{2}(\bm{y}^{\mathrm{ts}}_{i}(t),{\bm{{s}}}^{\mathrm{ts}})}{\sigma^{2}(\bm{y}^{\mathrm{ts}}_{i}(t))\sigma^{2}({\bm{{s}}}^{\mathrm{ts}})}.$ (12)
@@
As explained above, the QRP analyzes the variation of dynamics under different $s_{k}$ values. $R^{2}_{i}(t)$ should ideally be zero when no systematic variations dependent on the input $s_{k}$ are observed. However, in numerical simulations, $R^{2}_{i}(t)$ might accidentally exhibit nonzero values, sensitively reflecting the small deviations comparable to numerical errors. To avoid this numerical artifact, we introduce a threshold $\Delta_{\mathrm{th}}$ for the deviation of the operator across different instances. We employ the average absolute deviation $\Delta_{i}(t)$, defined as
@@
When $\Delta_{i}(t)$ is negligibly small, $R^{2}_{i}(t)$ calculated from the above $\langle O_{i}(t)\rangle$ should approach zero. We thus manually set $R^{2}_{i}(t)=0$ when $\Delta_{i}(t)<\Delta_{\mathrm{th}}$; we utilize $\Delta_{\mathrm{th}}=10^{-5}$, and $\Delta_{\mathrm{th}}$ dependency is presented in the Supplemental Information. This procedure does not qualitatively alter our methodology and is experimentally feasible.
@@
The data that support the findings of this study are available from the corresponding author upon reasonable request.
@@
The simulation codes used in this study are available from the corresponding author upon reasonable request.
@@ ACKNOWLEDGEMENTS
This research was supported by a Grant-in-Aid for Scientific Research on Innovative Areas “Quantum Liquid Crystals” (KAKENHI Grant No. JP19H05825) from JSPS of Japan and JST CREST (No. JP-MJCR18T2). K.K. was supported by the Program for Leading Graduate Schools (MERIT-WINGS) and JSPS KAKENHI Grant Number JP24KJ0872. The computation in this work has been done using the facilities of the Supercomputer Center, the Institute for Solid State Physics, the University of Tokyo.
@@
In the main text, we have demonstrated the efficacy of the quantum reservoir probing (QRP) in identifying quantum phase transitions through a statistical analysis of the effects of a local quantum quench, as described by the estimation performance $R^{2}_{i}(t)$ in Eq. (5) in the main text. Figures 2d and 3b in the main text illustrate the signatures of quantum phase transitions and quantum critical points, utilizing the mean value of $R^{2}_{i}(t)$, which is calculated over selected spatiotemporal indices $(i,t)$ to minimize extraneous effects from the edges of the systems. Particularly, in the case of the transverse-field Ising model in Fig. 2d, we have presented the field dependence of the mean estimation performance $\overline{R^{2}}$, which is obtained by averaging $R^{2}_{i}(t)$ for indices $(i,t)\in\Lambda\times\{0<t\leq 10\}$, where $\Lambda$ encompasses the central $11$ sites. Similarly for the ANNNI model in Fig. 3b, we have exhibited $\overline{R^{2}}$ defined over $(i,t)$ drawn from the set $\Lambda\times\{0<t\leq 15\}$, where $\Lambda$ also contains central $11$ sites. It is imperative to note that the quantum phases manifest distinguishable characteristics in $R^{2}_{i}(t)$ (Figs. 2a and 3a in the main text), and $\overline{R^{2}}$ serves as an ancillary tool to further clarify these characteristics. In other words, this averaging operation is implemented only to enhance the clarity of the results, and $\overline{R^{2}}$, defined as the simple average of local quantities, remains inherently local, and does not include nonlocal quantum effects. In this Supplementary Note, we present $\overline{R^{2}}$ defined for various subsets of spatiotemporal indices $(i,t)$, thus demonstrating the robustness of the QRP in detecting quantum phase transitions independent of the employed spatiotemporal subset.
@@
Figure S1: Average of the estimation performance in the transverse-field Ising model over various subsets of spatiotemporal indices. a The average $\overline{R^{2}}$ over subsets $(i,t)\in\Lambda\times\{0<t\leq 10\}$, where $\Lambda$ encompasses the central 11 sites (blue), central 7 sites (pink), central 15 sites (green), and central 21 sites (orange). b The average $\overline{R^{2}}$ over a subsets $(i,t)\in\Lambda\times\{t\}$, where $\Lambda$ encompasses the central 11 sites. The time window is $0<t\leq 10$ (blue), $2.5<t\leq 7.5$ (pink), $0<t\leq 7.5$ (green), and $5\leq t\leq 15$ (orange). The system size is $N=49$ in both (a) and (b).
@@
Figure S1 displays the average $\overline{R^{2}}$ in the transverse-field Ising model as a function of the field $g$ in the evolution Hamiltonian $\mathcal{H}(g)$. In Fig. S1a, we show $\overline{R^{2}}$ computed over various spatial index sets $\Lambda$, while the temporal index range is held constant at $\{0<t\leq 10\}$. Conversely, in Fig. S1b, we present $\overline{R^{2}}$ calculated over fixed sets $\Lambda$, which encompass the central eleven sites, while changing the temporal index range. While each curve corresponding to the subset $\{(i,t)\}$ exhibits distinct shapes, the qualitative characteristics are consistent: $\overline{R^{2}}$ decreases as it approaches $g=1.0$, reaches a minimum at $g_{c}$ due to the strong fluctuations, and subsequently rises with further increasing $g$. These patterns robustly indicate the occurrence of a quantum phase transition, with the quantum critical point at $g_{c}$ demarcating the transition between the ferromagnetically ordered phase and the quantum disordered phase.
@@
Figure S2 demonstrates the averaged estimation performance in the ANNNI model as a function of the field strength. Figure S2a illustrates $\overline{R^{2}}$ calculated over various spatial index sets $\Lambda$ with the temporal index range consistently set at $\{0<t\leq 15\}$. In addition, Fig. S2b illustrates $\overline{R^{2}}$ calculated over a fixed spacial indices set $\Lambda$ that includes the central eleven sites, while varying the temporal index range. Similar to the observations in the transverse-field Ising model, the qualitative behaviors of $\overline{R^{2}}$ in the ANNNI model remain consistent across different subsets $\{(i,t)\}$, with all results corroborating the occurrence of a quantum phase transition around $g_{c}\simeq 1.6$. However, a notable feature in the ANNNI model is the broader shape of the dip in $\overline{R^{2}}$ near the quantum critical point, which leads to variations in the identified field strength at which $\overline{R^{2}}$ reaches its minimum depending on the specific subset employed. Nevertheless, it is anticipated that this variance would decrease as the system size increases. A larger system would reduce the influence of edge effects, which are the principal cause of the subset $\{(i,t)\}$ dependency in $\overline{R^{2}}$. Consequently, in a larger system, the quantum critical point is expected to be identified with greater precision and clarity.
@@
Figure S2: Average of the estimation performance in the ANNNI model over various subsets of spatiotemporal indices. a The average $\overline{R^{2}}$ over subsets $(i,t)\in\Lambda\times\{0<t\leq 15\}$, where $\Lambda$ encompasses the central 11 sites (skyblue), central 7 sites (pink), central 15 sites (green), and central 21 sites (orange). b The average $\overline{R^{2}}$ over subsets $(i,t)\in\Lambda\times\{t\}$, where $\Lambda$ encompasses the central 11 sites. The time window is $0<t\leq 15$ (skyblue), $5\leq t\leq 10$ (pink), $2.5\leq t\leq 12.5$ (green), and $0<t\leq 20$ (orange). The system size is $N=49$ in both (a) and (b).
Supplementary Note 2: Dependence of the estimation performance on the deviation threshold
@@
As explained in the Methods section in the main text, we introduced the threshold for deviations in the dynamics over multiple cases in the evaluation of the estimation performance. The average absolute deviation $\Delta_{i}(t)$, defined in Eq. (13) in the main text, is employed as a criterion. $R^{2}_{i}(t)$ should be zero when $\Delta_{i}(t)$ is negligible at the level of the numerical error, however, it may accidentally become nonzero sensitively reflecting the small deviations. To circumvent such numerical artifacts, we manually set $R^{2}_{i}(t)=0$ when $\Delta_{i}(t)$ is smaller than the threshold $\Delta_{\mathrm{th}}$. Figures in the main text are obtained under $\Delta_{\mathrm{th}}=10^{-5}$, whereas the qualitative behaviors remain consistent with different $\Delta_{\mathrm{th}}$ as shown in the following.
@@
In Fig. S3, we present $R^{2}_{i}(t)$ in the transverse-field Ising model for three different $\Delta_{\mathrm{th}}$; the results with $\Delta_{\mathrm{th}}=10^{-5}$ are illustrated in Fig. 2a in the main text. In Fig. S3a, $R^{2}_{i}(t)$ without the threshold is presented, where $R^{2}_{i}(t)$ is depicted as obtained. Remarkably, we observe nonzero $R^{2}_{i}(t)$ even outside the light cone, likely due to numerical artifacts. In Figs. S3b and S3c, we introduce the threshold $\Delta_{\mathrm{th}}=10^{-6}$ and $\Delta_{\mathrm{th}}=10^{-4}$, respectively. The unphysical nonzero values of $R^{2}_{i}(t)$ outside the wavefronts are eliminated, mitigating numerical artifacts. Furthermore, the introduction of the threshold narrows the “width” of the wavefront, particularly evident at $g=0.5$. However, once the effects of the local quench have propagated (i.e., inside the wave fronts), the behavior of $R^{2}_{i}(t)$ changes minimally, and the phase-dependent behavior remains consistent. Therefore, the introduction of the threshold to the QRP does not qualitatively alter the outcomes.
@@
Figure S3: Dependence of the estimation performance on the deviation threshold in the transverse-field Ising model. The spatiotemporal representation of $R^{2}_{i}(t)$, as depicted in Fig. 2a in the main text. $R^{2}_{i}(t)$ is calculated with the threshold a $\Delta_{\mathrm{th}}=0.0$ (no threshold), b $\Delta_{\mathrm{th}}=10^{-6}$, and c $\Delta_{\mathrm{th}}=10^{-4}$.

### v3

@@ METHODS
To construct the initial state, we define $\mathcal{H}_{\mathrm{DMRG}}$ that yields $|\Psi_{k}^{\mathrm{in}}\rangle$ as its ground state obtained from the DMRG method [60, 61, 62]. The Hamiltonian $\mathcal{H}_{\mathrm{DMRG}}$ comprises the local magnetic fields, aligned with the product state specified in Eqs. (6) and (10), as well as the input magnetic field applied to the central spin to prepare it in the $s_{k}$-dependent state. The initial state is then obtained as the ground state of $\mathcal{H}_{\mathrm{DMRG}}$ using the DMRG method, expressed as a matrix product state with a bond dimension $\chi$. For the time-evolution $e^{-i\mathcal{H}t}|\Psi_{k}^{\mathrm{in}}\rangle$, we employ the time-dependent variational principle [63, 64, 65] with a bond dimension $\chi$ and a time step $\Delta t=0.005$. To maintain computational accuracy, the time evolution is conducted only up to the point before any anomalous behavior in the entanglement entropy is observed (refer to Supplementary Information). The tensor network calculations are implemented using the Julia version of the ITensor library [66, 67].
@@
From a uniform distribution $[0,1]$, a set of $256$ input values $\{s_{k}\}$ is sampled. Each input $s_{k}$ produces a distinct initial state $|\Psi_{k}^{\mathrm{in}}\rangle$, and the dynamics of operator $\langle O_{i}(t)\rangle_{k}$ is subsequently calculated. Among them, $l^{\mathrm{tr}}=128$ instances are designated for training, while the remaining $l^{\mathrm{ts}}=128$ instances are reserved for testing. Using a 2D vector $\mathbf{x}_{i,k}(t)=\left(\langle O_{i}(t)\rangle_{k},1\right)$, the linear transformation for estimating $s_{k}$ is expressed as $y_{i,k}(t)=\mathbf{x}_{i,k}(t)\mathbf{w}_{i}(t)$, where the weight vector is defined by $\mathbf{w}_{i}(t)=(w_{i,O}(t),w_{i,c}(t))^{\top}$.
In the training phase, an ($l^{\mathrm{tr}}\times 2$)-dimensional matrix $\mathrm{X}_{i}^{\mathrm{tr}}(t)=\{\mathbf{x}_{i,k}(t)\}_{k=1}^{l^{\mathrm{tr}}}$ is constructed by summing up the observed dynamics. The output, represented as an $l^{\mathrm{tr}}$-dimensional vector $\mathbf{y}_{i}^{\mathrm{tr}}(t)=\{y_{i,k}\}_{k=1}^{l^{\mathrm{tr}}}$, is calculated as $\mathbf{y}_{i}^{\mathrm{tr}}(t)=\mathrm{X}^{\mathrm{tr}}_{i}(t)\mathbf{w}_{i}(t)$. The weight vector is trained to closely match the output $\mathbf{y}_{i}^{\mathrm{tr}}(t)$ with the original input ${\mathbf{{s}}}^{\mathrm{tr}}=\{s_{k}\}_{k=1}^{l^{\mathrm{tr}}}$. The optimal solution, which minimizes the least squared error between these vectors, is given by
$\mathbf{w}_{i}(t)={\mathrm{X}_{i}^{\mathrm{tr}}}(t)^{+}{\mathbf{{s}}}^{\mathrm{tr}},$ (11)
where ${\mathrm{X}_{i}^{\mathrm{tr}}}(t)^{+}$ denotes the Moore-Penrose pseudoinverse-matrix of ${\mathrm{X}_{i}^{\mathrm{tr}}}(t)$.
In the testing phase, an ($l^{\mathrm{ts}}\times 2$)-dimensional matrix $\mathrm{X}_{i}^{\mathrm{ts}}(t)=\{\mathbf{x}_{i,k}(t)\}_{k=l^{\mathrm{tr}}+1}^{l^{\mathrm{tr}}+l^{\mathrm{ts}}}$ is similarly composed, yielding the output $\mathbf{y}_{i}^{\mathrm{ts}}(t)=\mathrm{X}_{i}^{\mathrm{ts}}(t)\mathbf{w}_{i}(t)$ using the trained weights. The estimation performance is assessed by comparing this testing output with the corresponding input ${\mathbf{{s}}}^{\mathrm{ts}}=\{s_{k}\}_{k=l^{\mathrm{tr}}+1}^{l^{\mathrm{tr}}+l^{\mathrm{ts}}}$ using the determination coefficient
$R^{2}_{i}(t)=\frac{\mathrm{cov}^{2}(\mathbf{y}^{\mathrm{ts}}_{i}(t),{\mathbf{{s}}}^{\mathrm{ts}})}{\sigma^{2}(\mathbf{y}^{\mathrm{ts}}_{i}(t))\sigma^{2}({\mathbf{{s}}}^{\mathrm{ts}})}.$ (12)
@@
As explained above, the QRP analyzes the variation of dynamics under different $s_{k}$ values. $R^{2}_{i}(t)$ should ideally be zero when no systematic variations dependent on the input $s_{k}$ are observed. However, in numerical simulations, $R^{2}_{i}(t)$ might accidentally exhibit nonzero values, sensitively reflecting the small deviations in $\langle O_{i}(t)\rangle_{k}$ comparable to numerical errors. To avoid this numerical artifact, we introduce a threshold $\Delta_{\mathrm{th}}$ for the deviation of the operator across different instances. We employ the average absolute deviation $\Delta_{i}(t)$, defined as
@@
When $\Delta_{i}(t)$ is negligibly small, $R^{2}_{i}(t)$ calculated from the above $\langle O_{i}(t)\rangle$ should approach zero. We thus manually set $R^{2}_{i}(t)=0$ when $\Delta_{i}(t)<\Delta_{\mathrm{th}}$; we utilize $\Delta_{\mathrm{th}}=10^{-5}$, and $\Delta_{\mathrm{th}}$ dependency is presented in the Supplementary Information. This procedure does not qualitatively alter our methodology and is experimentally feasible.
@@
The data that support the findings of this study are available from the corresponding author upon request.
@@
The simulation codes used in this study are available from the corresponding author upon request.
@@ ACKNOWLEDGEMENTS
K.K. thanks Frank Pollmann for insightful discussions. This research was supported by a Grant-in-Aid for Scientific Research on Innovative Areas “Quantum Liquid Crystals” (KAKENHI Grant No. JP19H05825) from JSPS of Japan and JST CREST (No. JP-MJCR18T2). K.K. was supported by the Program for Leading Graduate Schools (MERIT-WINGS), JSPS KAKENHI Grant Number JP24KJ0872, and JST BOOST Grant Number JPMJBS2418. The computation in this work has been done using the facilities of the Supercomputer Center, the Institute for Solid State Physics, the University of Tokyo.
@@
In the main text, we have demonstrated the efficacy of the quantum reservoir probing (QRP) in identifying quantum phase transitions through a statistical analysis of the effects of the local quantum quench, as described by the estimation performance $R^{2}_{i}(t)$ in Eq. (5) in the main text. Figures 2d and 3b in the main text illustrate the signatures of quantum phase transitions and quantum critical points utilizing the mean value of $R^{2}_{i}(t)$, which is calculated over selected spatiotemporal indices $(i,t)$ to minimize extraneous effects from the edges of the systems. Particularly, in the case of the transverse-field Ising model in Fig. 2d, we have presented the field dependence of the mean estimation performance $\overline{R^{2}}$, which is obtained by averaging $R^{2}_{i}(t)$ for indices $(i,t)\in\Lambda\times\{0<t\leq 7\}$, where $\Lambda$ encompasses the central nine sites. Similarly for the ANNNI model in Fig. 3b, we have exhibited $\overline{R^{2}}$ defined over $(i,t)$ drawn from the subset $\Lambda\times\{0<t\leq 5\}$, where $\Lambda$ also contains central seven sites. It is imperative to note that the quantum phases manifest distinguishable characteristics in the distribution of $R^{2}_{i}(t)$ (Figs. 2a and 3a in the main text), and $\overline{R^{2}}$ serves as an ancillary tool to further clarify these characteristics. In other words, this averaging operation is implemented only to enhance the clarity of the results, and $\overline{R^{2}}$, defined as the simple average of local quantities, remains inherently local and does not include nonlocal quantum effects. In this Supplementary Note, we present $\overline{R^{2}}$ defined for various subsets of spatiotemporal indices $(i,t)$ and demonstrate the robustness of the QRP in detecting quantum phase transitions independent of the employed spatiotemporal subset.
@@
Supplementary Fig. 1: Average of the estimation performance in the transverse-field Ising model over various subsets of spatiotemporal indices. a The average $\overline{R^{2}}$ over subsets $(i,t)\in\Lambda\times\{0<t\leq 7\}$, where $\Lambda$ encompasses the central 9 sites (blue), central 5 sites (pink), central 11 sites (green), and central 15 sites (orange). b The average $\overline{R^{2}}$ over subsets $(i,t)\in\Lambda\times\{t\}$, where $\Lambda$ encompasses the central nine sites. The time window is $0<t\leq 7$ (blue), $3\leq t\leq 7$ (pink), $0<t\leq 4$ (green), and $2\leq t\leq 5$ (orange). The system size is $N=49$ in both (a) and (b).
@@
Supplementary Figure 1 displays the average $\overline{R^{2}}$ in the transverse-field Ising model as a function of the field $g$ in the Hamiltonian $\mathcal{H}(g)$. In Supplementary Fig. 1a, we show $\overline{R^{2}}$ computed over various spatial index sets $\Lambda$, while the temporal index range is held constant at $\{0<t\leq 7\}$. Conversely, in Supplementary Fig. 1b, we present $\overline{R^{2}}$ calculated over the fixed set $\Lambda$ that encompasses the central nine sites, while changing the temporal index range. Although each curve corresponding to a different subset $\{(i,t)\}$ exhibits a distinct shape, the qualitative characteristics are consistent: $\overline{R^{2}}$ decreases as $g$ approaches $g_{c}=1.0$, reaches a minimum at $g_{c}$ due to the strong fluctuations, and subsequently rises with further increasing $g$. These patterns robustly indicate the occurrence of the quantum phase transition at the quantum critical point $g_{c}$, regardless of the averaging spatiotemporal region.
@@
Supplementary Figure 2 shows the averaged estimation performance in the ANNNI model as a function of the field strength. Supplementary Figure 2a illustrates $\overline{R^{2}}$ calculated over various spatial index sets $\Lambda$ with the temporal index range consistently set at $\{0<t\leq 5.2\}$. In addition, Supplementary Fig. 2b illustrates $\overline{R^{2}}$ calculated over a fixed spatial indices set $\Lambda$ that includes the central five sites, while varying the temporal index range. Similar to the observations in the transverse-field Ising model, the qualitative behaviors of $\overline{R^{2}}$ in the ANNNI model remain consistent across different subsets $\{(i,t)\}$, with all results corroborating the occurrence of the quantum phase transition around $g_{c}\simeq 1.7$. However, a notable feature in the ANNNI model is the broader shape of the dip in $\overline{R^{2}}$ near the quantum critical point, which leads to variations in the identified field strength at which $\overline{R^{2}}$ reaches its minimum, depending on the specific subset employed. Nevertheless, with increasing the system size, it is anticipated that this variance would decrease and the quantum critical point would be identified with greater precision and clarity.
@@
Supplementary Fig. 2: Average of the estimation performance in the ANNNI model over various subsets of spatiotemporal indices. a The average $\overline{R^{2}}$ over subsets $(i,t)\in\Lambda\times\{0<t\leq 5\}$, where $\Lambda$ encompasses the central 7 sites (skyblue), central 3 sites (pink), central 11 sites (green), and central 13 sites (orange). b The average $\overline{R^{2}}$ over subsets $(i,t)\in\Lambda\times\{t\}$, where $\Lambda$ encompasses the central 7 sites. The time window is $0<t\leq 5$ (skyblue), $3\leq t\leq 5$ (pink), $0<t\leq 3$ (green), and $1\leq t\leq 4$ (orange). The system size is $N=49$ in both (a) and (b).
Supplementary Note 2: Dynamics of the entanglement entropy
In the time evolution of a quantum system, the growth of entanglement entropy is a crucial aspect to monitor. Specifically, entanglement increases rapidly and spreads throughout the system, eventually exceeding the computational capacity to accurately represent the quantum state beyond a certain temporal threshold. This limitation imposes significant constraints on the achievable simulation timescale. To explore this, we examine the entanglement entropy between two subsystems, $A$ and $B$, defined as $S=-\mathrm{Tr}_{A}(\rho_{A}\ln\rho_{A})$, where $\rho_{A}$ denotes the reduced density matrix $\rho_{A}=\mathrm{Tr}_{B}\rho$. Since a matrix product state with bond dimension $\chi$ can only capture the entanglement entropy up to $S\leq\ln\chi$, a temporal threshold arises for precise dynamical calculations, beyond which the state cannot be accurately approximated due to increased entanglement. This typically leads to anomalous behavior in the dynamics of the entanglement entropy; for instance, the growth in $S$ deviates from the expected behavior extrapolated from early times, such as reaching saturation near a value of $\ln\chi$. To ensure numerical accuracy, it is essential to determine the maximum simulation time by closely monitoring this entanglement evolution.
In Supplementary Fig. 3, we present the half-chain entanglement entropy for each model investigated in the main text. The initial state $|\Psi_{k}^{\mathrm{in}}\rangle$ in Eq. (6) or (10) is prepared with a representative input value of $s_{k}=0.5$. For all models and parameter settings, the entanglement entropy generally shows linear growth during early times, with oscillatory behavior in certain instances. At a specific time point (indicated by the gray dashed line in Supplementary Fig. 3), the linear growth rate begins to deviate due to the constraint of maximal entanglement entropy. We consider the dynamics beyond this point to be potentially unreliable and focus only on the data prior to the onset of this anomalous behavior. Importantly, as demonstrated in the main text, the signatures of the quantum phase transitions become evident as early as $t\sim 1$ or even earlier, which can be captured with high precision both numerically and experimentally.
Supplementary Fig. 3: Dynamics of the entanglement entropy. The time evolution of the half-chain entanglement entropy $S$ examined for (a) the transverse-field Ising model, (b) the ANNNI model, (c) the cluster model, and (d) the cluster model in a magnetic field. Each line corresponds to a distinct model parameter. The initial state is prepared with a typical input value $s_{k}=0.5$. The system size is $N=49$, with a maximum bond dimension of $\chi=256$. The gray dashed line denotes the maximum time utilized to generate the results shown in the main text for each model.
Supplementary Note 3: Dependence of the estimation performance on the deviation threshold
@@
As explained in the Methods section in the main text, we introduced the threshold for deviations in the dynamics over multiple instances in the evaluation of the estimation performance. The average absolute deviation $\Delta_{i}(t)$, defined in Eq. (13) in the main text, is employed as a criterion. $R^{2}_{i}(t)$ should be zero when $\Delta_{i}(t)$ is negligible at the level of the numerical error, however, it may accidentally become nonzero, sensitively reflecting the small deviations. To circumvent such numerical artifacts, we manually set $R^{2}_{i}(t)=0$ when $\Delta_{i}(t)$ is smaller than the threshold $\Delta_{\mathrm{th}}$. Figures in the main text are obtained under $\Delta_{\mathrm{th}}=10^{-5}$, whereas the qualitative behaviors remain consistent with different $\Delta_{\mathrm{th}}$ as shown in the following.
@@
In Supplementary Fig. 4, we present $R^{2}_{i}(t)$ in the transverse-field Ising model for three different $\Delta_{\mathrm{th}}$; the results with $\Delta_{\mathrm{th}}=10^{-5}$ are illustrated in Fig. 2a in the main text. In Supplementary Fig. 4a, $R^{2}_{i}(t)$ without the threshold is presented, where $R^{2}_{i}(t)$ is depicted as obtained. We observe nonzero $R^{2}_{i}(t)$ even outside the light cone, likely due to numerical artifacts. In Supplementary Fig. 4b and 4c, we introduce the threshold $\Delta_{\mathrm{th}}=10^{-6}$ and $\Delta_{\mathrm{th}}=10^{-4}$, respectively. The introduction of the threshold narrows the “width” of the wavefront in the early stages, and outside the wavefronts the unphysical nonzero values of $R^{2}_{i}(t)$ are eliminated due to reduced numerical sensitivity. Once the effects of the local quench have propagated (i.e., inside the wave fronts), the behavior of $R^{2}_{i}(t)$ remains largely unchanged, and the phase-dependent behavior remains consistent. Therefore, the introduction of the threshold to the QRP does not fundamentally alter the outcomes of the QRP.
@@
Supplementary Fig. 4: Dependence of the estimation performance on the deviation threshold in the transverse-field Ising model. The spatiotemporal representation of $R^{2}_{i}(t)$, as depicted in Fig. 2a in the main text. $R^{2}_{i}(t)$ is calculated with the threshold a $\Delta_{\mathrm{th}}=0.0$ (no threshold), b $\Delta_{\mathrm{th}}=10^{-6}$, and c $\Delta_{\mathrm{th}}=10^{-4}$.